# Diffusion Models, Image Super-Resolution And Everything: A Survey

### 01 Jan 2024 v1

@@ 3.2 Denoising Diffusion Probabilistic Models (DDPMs)
Ho et al. [6] recommend reweighting the VLB for improved sample quality:
$\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(1,T)\\ \mathbf{z}_{0}\sim q(\mathbf{z}_{0})\\ \epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)\end{subarray}}\left[\lambda(t)\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t)\|^{2}\right],$ (19)
where $\lambda(t)>0$ is a weighting function, $\mathbf{z}_{t}$ is derived by Equation 14 and $\epsilon_{\theta}$ is a NN that predicts the noise vector $\epsilon$ given $\mathbf{z}_{t}$ and $t$. For a specific choice of $\lambda(t)$, Equation 19 reduces to Equation 18. We present this formulation to show later that it intriguingly bears a resemblance to the objective of denoising score matching for training Score-based Generative Models (SGMs), which will be introduced next, and draws the connection between both approaches.
3.3 Score-based Generative Models (SGMs)
@@
Backward Diffusion: To revert the noise during the backward diffusion, we need to approximate $\nabla_{\mathbf{z}_{t}}\log{q(\mathbf{z}_{t})}$ and choose a method for estimating the intermediate states $\mathbf{z}_{t}$ from that approximation. For the gradient approximation at each time step $t$, we use a trained predictor, denoted as $s_{\theta}$ and called Noise-Conditional Score Network (NCSN), such that $s_{\theta}(\mathbf{z}_{t},t)\approx\nabla_{\mathbf{z}_{t}}\log{q(\mathbf{z}_{t})}$ [9]. The training of the NCSN will be covered in the next section; for now, we focus on the sampling process using NCSN. Sampling with NCSN involves generating the intermediate states $\mathbf{z}_{t}$ through an iterative approach, using $s_{\theta}(\mathbf{z}_{t},t)$. Note that this iterative process is different from the iterations done during the diffusion as it addresses solely the generation of $\mathbf{z}_{t}$.
@@
As mentioned in the chapter before, Equation 19 resembles Equation 26, which becomes clear once we set $\epsilon_{\theta}(\mathbf{z}_{t},t)=-\sigma_{t}s_{\theta}(\mathbf{z}_{t},t)$ in Equation 19. Thus, there is a mathematical connection between DDPMs and SGMs.
3.4 Stochastic Differential Equations (SDEs)
So far, we have discussed DMs that deal with finite time steps. A generalization to infinite time steps is made by formulating these as solutions to Stochastic Differential Equations (SDEs), also known as Score SDEs [51].
@@ 3.4 Stochastic Differential Equations (SDEs)
where f and g are the drift and diffusion functions, respectively, and $\mathbf{w}$ is the standard Wiener process (also known as Brownian motion). The SDE for DDPMs is given by:
@@
3.5 Relation to other Image SR Generative Models
The exploration of generative models in image SR is fascinating because it offers the potential to improve the quality of upsampled images beyond interpolation significantly. They differ primarily in how they approach the task of generating HR images from LR inputs. These differences stem from the underlying architecture and training objectives. While they offer significant advantages, they come with their individual set of challenges, such as training stability and computational costs.
@@ 3.5 Relation to other Image SR Generative Models
4 Improvements for Diffusion Models
@@ 4.1 Efficient Sampling
Lastly, truncated diffusion is a technique that improves speed by prematurely ending the forward diffusion process [65, 66]. This early termination results in outputs that are not purely Gaussian random noise, presenting computational challenges. These challenges are overcome using proxy distributions from pre-trained VAEs or GANs, which match the diffused data distribution and facilitate efficient backward diffusion.
@@ 4.2 Improved Likelihood
5 Diffusion Models for Image SR
So far, our discussion has revolved around the theoretical framework of DMs. This section explores the practical applications and recent publications in image SR.
5.1 Concrete Realization of DDPMs
@@ 5.2 Alternative Domains
5.2.1 Latent Space Domain
Models like SR3 [42] and SRDiff [74] have achieved high-quality SR results by operating in the pixel domain. However, these models are computationally intensive due to their iterative nature and the high-dimensional calculations in RGB space. To reduce computational demands, Latent Diffusion Models (LDMs), introduced by Rombach et al. [75], perform diffusion processes in a low-dimensional latent space. This approach significantly lowers resource requirements without compromising performance and allows the reuse of latent representations across multiple models.
LDMs typically follow a two-stage process. First, a pre-trained autoencoder, like the VQ-GAN [11], compresses input data into a low-dimensional latent representation. This encoder is trained using perceptual and adversarial losses. In the second stage, a time-conditional U-Net model is trained in the latent space, incorporating image-specific biases. For image SR, low-resolution images are concatenated as a condition to this model.
@@ 5.2.1 Latent Space Domain
TABLE I: PSNR and SSIM comparison on CelebA-HQ face SR 16$\times$16 $\rightarrow$ 128$\times$128. Consistency measures MSE (×$\times 10^{-5}$) between LR inputs and the downsampled SR outputs.
Methods PSNR $\uparrow$ SSIM $\uparrow$ Consistency $\downarrow$
PULSE [81] 16.88 0.44 161.1
FSRGAN [82] 23.01 0.62 33.8
SR3 (regression) [42] 23.96 0.69 2.71
SR3 (diffusion) [42] 23.04 0.65 2.68
DiWa [83] 23.34 0.67 -
IDM [73] 24.01 0.71 2.14
5.2.2 Wavelet Target Domain
@@
5.2.3 Residual Target Domain
As mentioned, SRDiff [74] is the pioneering work that advocates for shifting the image generation process into the residual space. This strategy aims to generate the difference between the upsampled LR and the HR image. This enables the DM to focus on residual details, speeds up convergence, and stabilizes the training [14]. Another notable contribution is ResDiff [88], which also seeks to predict the residuals. Whang et al. [89] also employs residual predictions as a fundamental component of their predict-and-refine approach for image deblurring. In a different vein, Yue et al. [90] presents a technique that constructs a Markov chain of transformations between HR and LR images by manipulating the residual between them. This novel approach substantially enhances transition efficiency, providing an innovative perspective on image manipulation strategies.
TABLE II: Results for 4× SR of general images on DIV2K val. Note that EDSR, FxSR-PD, CAR, and RRDB are regression-based methods that generally produce better PSNR and SSIM scores than generative approaches [42].
@@ 5.2.3 Residual Target Domain
5.3 Conditioning Diffusion Models
DMs heavily depend on conditioning information to guide the sampling process toward a reasonable HR prediction. This section reviews various methods for seamlessly integrating conditioning information into backward diffusion.
5.3.1 Low Resolution Reference
5.3.2 Super-Resolved Reference
@@
5.3.3 Feature Reference
Another avenue for conditioning involves relevant features extracted from pre-trained networks. SRDiff [74] leverages a pre-trained encoder to encode LR image features at each step of the backward diffusion. These features serve as guidance, aiding in the generation of higher-resolution outputs. Implicit DMs (IDMs) [73] take a different approach by conditioning their denoising network with a neural representation, which enables the learning of a continuous representation at various scales. They encode the image as a function within continuous space and seamlessly integrate it into the DM. These extracted features are adapted to multiple scales and are used across multiple layers within the DMs. To comprehensively understand the performance differences between these approaches, comparisons can be found in Table I and Table II.
5.3.4 Exploiting Text-To-Image
@@
5.4 Guidance in Training
We have examined the learning of conditional distributions, which form the backbone of image SR [108, 42]. As such, the condition $\mathbf{x}$, such as an LR image, is integrated into the backward diffusion, i.e., $p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{x}\right)$ for DDPMs or in $s_{\theta}(\mathbf{z}_{t},\mathbf{x},t)$ for SGMs/SDEs. However, this simple formulation can result in a model that overlooks the conditioning information.
@@ 5.4.1 Classifier Guidance
5.4.2 Classifier-Free Guidance
@@ 5.6 Corruption Space
HR
SR3
@@
5.7 Color Shifting
As a result of high computational costs, DMs can occasionally suffer from color shifting when limited hardware necessitates smaller batch sizes or shorter learning periods [132]. An example with SR3 is shown in Figure 5.
As presented by StableSR, a straightforward modification can address this issue by performing color normalization by adjusting the mean and variance with those of the LR image on the generated image [99]. Mathematically, it gives the following equation:
@@ 5.7 Color Shifting
where $c\in\{r,g,b\}$ denotes the color channel, and $\mu_{\mathbf{z}_{0}}^{c}$ and $\sigma^{c}_{\mathbf{\mathbf{z}_{0}}}$ (or $\mu_{\mathbf{x}}^{c}$ and $\sigma_{\mathbf{x}}^{c}$) are the mean and standard variance from the $c$-th channel of the predicted image $\mathbf{z}_{0}$ (or the input image $\mathbf{x}$), respectively.
You Only Diffuse Areas (YODA) [121], which targets diffusion on important image areas more frequently through time-dependent masks generated with DINO [122], also mitigates the color shift effect for image SR. This suggests that properly defined architecture and diffusion design are crucial to omit this effect. Further analysis of why this effect emerges must be obtained in future work.
@@
5.8 Architecture Designs for Denoising
The design of the denoising model in DMs offers a range of options. The majority of DMs adopt the use of U-Net, as noted in most literature [110]. SR3 [42], for instance, employs residual blocks from BigGAN [110] and re-scales skip connections by a factor of $\frac{1}{\sqrt{2}}$. SRDiff takes a similar approach [74], although it opts for vanilla residual blocks without the re-scaling of skip connections and uses a LR encoder to incorporate the information of the LR image during the backward diffusion.
Whang et al. [89] exploit an initial predictor to combine the strengths of deterministic image SR models and DMs, as shown in Figure 6. It has the advantage that the DM only needs to learn the residuals that the deterministic image SR model (initial predictor) fails to predict, hence simplifying the learning target. Additionally, the removal of self-attention, positional encodings and group normalization from the SR3 U-Net enables their model to support arbitrary resolutions.
An initial predictor is also employed in the wavelet-based approach DiWa [83]. Moreover, wavelet SR models, such as DWSR [85] – a simple sequence of convolution layers of depth 10 – are utilized for denoising prediction in the wavelet domain. In WaveDM [86], a deterministic U-Net predictor is used for the high-frequency band, while diffusion is applied in the low-frequency band.
6 Domain-Specific Applications
@@ 6.1 Medical Imaging
Magnetic Resonance Imaging (MRI) scans are widely used to aid patient diagnosis but can often be of low quality and corrupted with noise. Chung et al. [136] propose a combined denoising and SR network referred to as R2222D2+limit-from22+2 + (Regularized Reverse Diffusion Denoiser $+$ SR). They perform denoising of the MRI scans, followed by an SR module. Inspired by CCDF (i.e., a zero-shot method) from Chung et al. [123], they start their backward diffusion from an initial noisy image instead of pure Gaussian noise. The reverse SDE is solved using a non-parametric, eigenvalue-based method. In addition, they restrict the stochasticity of the DMs through low-frequency regularization. Particularly, they maintain low-frequency information while correcting the high-frequency ones to produce sharp and super-resolved MRI scans.
Mao et al. [137] addresses the lack of diffusion-based multi-contrast MRI SR methods. They propose a Disentangled Conditional Diffusion model (DisC-Diff) to leverage a multi-conditional fusion strategy based on representation disentanglement, enabling high-quality HR image sampling. Specifically, they employ a disentangled U-Net with multiple encoders to extract latent representations and use a novel joint disentanglement and Charbonnier loss function to learn representations across MRI contrasts. They also implement curriculum learning and improve their MRI model for varying anatomical complexity by gradually increasing the difficulty of training images.
6.2 Blind Face Restoration

### 06 Feb 2024 v2

@@ III-B Denoising Diffusion Probabilistic Models (DDPMs)
III-C Score-based Generative Models (SGMs)
@@
Backward Diffusion: To revert the noise during the backward diffusion, we need to approximate $\nabla_{\mathbf{z}_{t}}\log{q(\mathbf{z}_{t})}$ and choose a method for estimating the intermediate states $\mathbf{z}_{t}$ from that approximation. For the gradient approximation at each time step $t$, we use a trained predictor, denoted as $s_{\theta}$ and called Noise-Conditional Score Network (NCSN), such that $s_{\theta}(\mathbf{z}_{t},t)\approx\nabla_{\mathbf{z}_{t}}\log{q(\mathbf{z}_{t})}$ [7].
The training of the NCSN will be covered in the next section; for now, we focus on the sampling process using NCSN. Sampling with NCSN involves generating the intermediate states $\mathbf{z}_{t}$ through an iterative approach, using $s_{\theta}(\mathbf{z}_{t},t)$. Note that this iterative process is different from the iterations done during the diffusion as it addresses solely the generation of $\mathbf{z}_{t}$. This is a key difference to DDPMs as $\mathbf{z}_{t}$ needs to be sampled iteratively, whereas DDPMs directly predict $\mathbf{z}_{t}$ from $\mathbf{z}_{t+1}$.
@@
III-D Stochastic Differential Equations (SDEs)
So far, we have discussed DMs that deal with finite time steps. A generalization to infinite time steps is made by formulating these as solutions to Stochastic Differential Equations (SDEs), also known as Score SDEs [17]. In fact, we can view SGMs and DDPMs as discretizations of a continuous-time SDE. SDEs are not entirely bound to diffusion models, as they are a mathematical concept describing stochastic processes. As such, they fit perfectly to describe the processes we want to simulate in diffusion models. Like previously, data is perturbed in a general diffusion process but generalized to an infinite number of noise scales.
@@ III-D Stochastic Differential Equations (SDEs)
where f and g are the drift and diffusion functions, respectively, and $\mathbf{w}$ is the standard Wiener process (also known as Brownian motion). One benefit of this generalized formulation is that DDPMs and SGMs can be formulated uniformly in this setting. The SDE for DDPMs is given by:
@@
Figure 2: Conceptual overview of generative models (GANs, VAEs, NFs, and DMs).
III-E Relation between Diffusion Models
As highlighted in the SDE section, we can describe both variations, namely SGMs and DDPMs, with SDEs. We can also showcase this close relationship by reformulating the optimization targets. For DDPMs, we saw in Equation 18 that
$\text{KL}(q\left(\mathbf{z}_{0},...,\mathbf{z}_{T}\right)\|p_{\theta}\left(\mathbf{z}_{0},...,\mathbf{z}_{T}\right))$
$\stackrel{{\scriptstyle(ii)}}{{\geq}}\mathbb{E}\left[-\log p_{\theta}\left(\mathbf{z}_{0}\right)\right]+c$
is minimized. By reweighting the VLB, as Ho et al. [4] recommends for improved sample quality, we can further derive:
$\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(1,T)\\ \mathbf{z}_{0}\sim q(\mathbf{z}_{0})\\ \epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)\end{subarray}}\left[\lambda(t)\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t)\|^{2}\right],$
where $\lambda(t)>0$ is a weighting function. If we now take the optimization target in Equation 25 of SGMs, which was
$\mathop{\mathbb{E}}_{\begin{subarray}{c}t\sim\mathcal{U}(1,T)\\ \mathbf{z}_{0}\sim q(\mathbf{z}_{0})\\ \epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)\end{subarray}}\left[\lambda(t)\|\epsilon+\sigma_{t}s_{\theta}(\mathbf{z}_{t},t)\|^{2}\right]+c,$
the relationship between DDPMs and SGMs becomes clear once we set $\epsilon_{\theta}(\mathbf{z}_{t},t)=-\sigma_{t}s_{\theta}(\mathbf{z}_{t},t)$. As the constant $c$ becomes irrelevant for the optimization, we can see once again that there is a mathematical connection between DDPMs and SGMs.
III-F Relation to other Image SR Generative Models
The exploration of generative models in image SR is fascinating because it offers the potential to improve the quality of upsampled images beyond interpolation significantly. They differ primarily in how they approach the task of generating HR images from LR inputs and are illustrated in Figure 2. These differences stem from the underlying architecture and training objectives. While they offer significant advantages, they come with their individual set of challenges, such as training stability and computational costs.
@@ III-F Relation to other Image SR Generative Models
IV Improvements for Diffusion Models
@@ IV-A Efficient Sampling
Lastly, truncated diffusion is a technique that improves speed by prematurely ending the forward diffusion process [67, 68]. This early termination results in outputs that are not purely Gaussian noise, presenting computational challenges. These challenges are overcome using proxy distributions from pre-trained VAEs or GANs, which match the diffused data distribution and facilitate efficient backward diffusion.
@@ IV-B Improved Likelihood
V Diffusion Models for Image SR
So far, our discussion has revolved around the theoretical framework of DMs. This section explores the practical applications and recent publications in image SR. We will discuss concrete realizations of diffusion models, which are predominantly DDPMs. Next, we will introduce alternative input domains for DDPMs and strategies to incorporate various ways of conditioning information (LR image) and how to improve conditioning with guidance. Moreover, we will explore zero-shot approaches, as well as for SR specifically interesting research areas, such as corruption spaces, color-shifting and architectural designs.
V-A Concrete Realization of Diffusion Models
While SGMs provide considerable flexibility in design, the prevailing trend within the field of image leans towards the implementation of DDPMs. Firstly, DDPMs are characterized by their straightforward implementation, which reduces the barrier to entry for researchers and developers alike. This ease of implementation is a significant advantage, as it allows for quicker development cycles and easier replication of results. Additionally, the inherent flexibility offered by SGMs, while advantageous in creating customized solutions, introduces complexity due to the multitude of design variables that need to be considered. This complexity can pose a challenge in research settings, where rigorously evaluating the impact of each variable (e.g., different sampling algorithms) becomes a cumbersome process. Moreover, the growing body of literature featuring DDPM implementations has contributed to their increasing popularity within the community. As more studies adopt and validate DDPMs for image SR applications, a virtuous cycle is created, where the familiarity and proven effectiveness of DDPMs encourage further adoption. This trend towards DDPMs is further reinforced by cumulative knowledge and shared experiences, making them a more common choice for image SR.
@@ V-B Alternative Domains
V-B1 Latent Space Domain
Models like SR3 [15] and SRDiff [76] have achieved high-quality SR results by operating in the pixel domain. However, these models are computationally intensive due to their iterative nature and the high-dimensional calculations in RGB space. To reduce computational demands, Latent Diffusion Models (LDMs), introduced by Rombach et al. [13], perform diffusion processes in a low-dimensional latent space. This approach significantly lowers resource requirements without compromising performance and allows the reuse of latent representations across multiple models. LDMs typically follow a two-stage process. First, a pre-trained autoencoder, like the VQ-GAN [9], compresses input data into a low-dimensional latent representation. This encoder is trained using perceptual and adversarial losses. In the second stage, a time-conditional U-Net model is trained in the latent space, incorporating image-specific biases. For image SR, low-resolution images are concatenated as a condition.
@@ V-B1 Latent Space Domain
V-B2 Wavelet Target Domain
@@
Figure 4: Overview of different ways to condition a diffusion-based SR model. One can incorporate the LR image, the SR image, its features, or a text-embedding of the LR image during the backward diffusion. Note that also combinations of different conditioning methods are possible.
V-B3 Residual Target Domain
As mentioned, SRDiff [76] is the pioneering work that advocates for shifting the image generation process into the residual space. This strategy aims to generate the difference between the upsampled LR and the HR image. This enables the DM to focus on residual details, speeds up convergence, and stabilizes the training [16]. Another notable contribution is ResDiff [87], which also seeks to predict the residuals. Whang et al. [88] also employs residual predictions as a fundamental component of their predict-and-refine approach for image deblurring.
In a different vein, Yue et al. [89] presents a technique that constructs a Markov chain of transformations between HR and LR images by manipulating the residual between them. This novel approach substantially enhances transition efficiency, providing an innovative perspective on image manipulation strategies.
V-C Conditioning Diffusion Models
DMs heavily depend on conditioning information to guide the sampling process toward a reasonable HR prediction. One common strategy is to incorporate the LR image during the backward diffusion. However, there are also other strategies that vary from the standard protocol. This section reviews various methods (see Figure 4) for seamlessly integrating conditioning information into backward diffusion.
V-C1 Low Resolution Reference
V-C2 Super-Resolved Reference
TABLE I: Results for 4× SR of general images on DIV2K val. Note that EDSR, FxSR-PD, CAR, and RRDB are regression-based methods that generally produce better PSNR and SSIM scores than generative approaches [15].
@@ V-C2 Super-Resolved Reference
TABLE II: PSNR and SSIM comparison on CelebA-HQ face SR 16$\times$16 $\rightarrow$ 128$\times$128. Consistency measures MSE (×$\times 10^{-5}$) between LR inputs and the downsampled SR outputs.
Methods PSNR $\uparrow$ SSIM $\uparrow$ Consistency $\downarrow$
PULSE [97] 16.88 0.44 161.1
FSRGAN [98] 23.01 0.62 33.8
SR3 (regression) [15] 23.96 0.69 2.71
SR3 (diffusion) [15] 23.04 0.65 2.68
DiWa [84] 23.34 0.67 -
IDM [75] 24.01 0.71 2.14
@@
V-C3 Feature Reference
Another avenue for conditioning involves relevant features extracted from pre-trained networks. SRDiff [76] leverages a pre-trained encoder to encode LR image features at each step of the backward diffusion. These features serve as guidance, aiding in the generation of higher-resolution outputs. Implicit DMs (IDMs) [75] take a different approach by conditioning their denoising network with a neural representation, which enables the learning of a continuous representation at various scales. They encode the image as a function within continuous space and seamlessly integrate it into the DM. These extracted features are adapted to multiple scales and are used across multiple layers within the DMs. To comprehensively understand the performance differences between these approaches, comparisons can be found in Table II and Table I.
V-C4 Exploiting Text-To-Image
@@
V-D Guidance in Training
We have examined the learning of conditional distributions, which form the backbone of image SR [109, 15]. As such, the condition $\mathbf{x}$, such as an LR image, is integrated into the backward diffusion, i.e., $p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{x}\right)$ for DDPMs or in $s_{\theta}(\mathbf{z}_{t},\mathbf{x},t)$ for SGMs/SDEs. However, this simple formulation can result in a model that overlooks the conditioning.
@@ V-D1 Classifier Guidance
V-D2 Classifier-Free Guidance
@@ V-F Corruption Space
Another avenue for alternative corruption space is presented by Inversion by Direct Iteration (InDI) [132]. InDI delineates a direct mapping strategy, efficiently bridging the gap between the two quality spaces without the iterative refinement typically required by conventional diffusion processes. The intrinsic flexibility and the direct mapping capability of InDI propose intriguing possibilities for enhancing image quality, suggesting a potent avenue for research exploration. The potential integration of InDI’s principles with those of conditional diffusion models could offer substantial advancements in the field of image SR. It is anticipated that a detailed examination and discussion of InDI within the broader scope of diffusion-based image enhancement could yield valuable insights and contribute significantly to the ongoing development of generative models in image processing.
@@
V-G Color Shifting
As a result of high computational costs, DMs can occasionally suffer from color shifting when limited hardware necessitates smaller batch sizes or shorter learning periods [133]. An example with SR3 is shown in Figure 8. As presented by StableSR, a straightforward modification can address this issue by performing color normalization by adjusting the mean and variance with those of the LR image on the generated image [100]. Mathematically, it gives the following equation:
@@ V-G Color Shifting
where $c\in\{r,g,b\}$ denotes the color channel, and $\mu_{\mathbf{z}_{0}}^{c}$ and $\sigma^{c}_{\mathbf{\mathbf{z}_{0}}}$ (or $\mu_{\mathbf{x}}^{c}$ and $\sigma_{\mathbf{x}}^{c}$) are the mean and standard variance from the $c$-th channel of the predicted image $\mathbf{z}_{0}$ (or the input image $\mathbf{x}$), respectively. You Only Diffuse Areas (YODA) [121], which targets diffusion on important image areas more frequently through time-dependent masks generated with DINO [122], also mitigates the color shift effect for image SR. This suggests that properly defined architecture and diffusion design are crucial to omit this effect. Further analysis of why this effect emerges must be obtained in future work.
V-H Architecture Designs for Denoising
The design of the denoising model in DMs offers a range of options. The majority of DMs adopt the use of U-Net, as noted in most literature [110]. SR3 [15], for instance, employs residual blocks from BigGAN [110] and re-scales skip connections by a factor of $\frac{1}{\sqrt{2}}$. SRDiff takes a similar approach [76], although it opts for vanilla residual blocks without the re-scaling of skip connections and uses a LR encoder to incorporate the information of the LR image during the backward diffusion. Whang et al. [88] exploit an initial predictor to combine the strengths of deterministic image SR models and DMs, as shown in Figure 9. It has the advantage that the DM only needs to learn the residuals that the deterministic image SR model (initial predictor) fails to predict, hence simplifying the learning target. Additionally, the removal of self-attention, positional encodings and group normalization from the SR3 U-Net enables their model to support arbitrary resolutions. An initial predictor is also employed in the wavelet-based approach DiWa [84]. Moreover, wavelet SR models, such as DWSR [83] – a simple sequence of convolution layers of depth 10 – are utilized for denoising prediction in the wavelet domain. In WaveDM [85], a deterministic U-Net predictor is used for the high-frequency band, while diffusion is applied in the low-frequency band.
@@ V-H Architecture Designs for Denoising
VI Domain-Specific Applications
@@ VI-A Medical Imaging
Magnetic Resonance Imaging (MRI) scans are widely used to aid patient diagnosis but can often be of low quality and corrupted with noise. Chung et al. [137] propose a combined denoising and SR network referred to as R2222D2+limit-from22+2 + (Regularized Reverse Diffusion Denoiser $+$ SR). They perform denoising of the MRI scans, followed by an SR module. Inspired by CCDF (i.e., a zero-shot method) from Chung et al. [123], they start their backward diffusion from an initial noisy image instead of pure Gaussian noise. The reverse SDE is solved using a non-parametric, eigenvalue-based method. In addition, they restrict the stochasticity of the DMs through low-frequency regularization. Particularly, they maintain low-frequency information while correcting the high-frequency ones to produce sharp and super-resolved MRI scans. Mao et al. [138] addresses the lack of diffusion-based multi-contrast MRI SR methods. They propose a Disentangled Conditional Diffusion model (DisC-Diff) to leverage a multi-conditional fusion strategy based on representation disentanglement, enabling high-quality HR image sampling. Specifically, they employ a disentangled U-Net with multiple encoders to extract latent representations and use a novel joint disentanglement and Charbonnier loss function to learn representations across MRI contrasts. They also implement curriculum learning and improve their MRI model for varying anatomical complexity by gradually increasing the difficulty of training images.
VI-B Blind Face Restoration